### v1

@@ Abstract
Reinforcement Learning (RL) agents are increasingly used to simulate sophisticated cyberattacks, but their decision-making processes remain opaque, hindering trust, debugging, and defensive preparedness. In high-stakes cybersecurity contexts, explainability is essential for understanding how adversarial strategies are formed and evolve over time. In this paper, we propose a unified, multi-layer explainability framework for RL-based attacker agents that reveals both strategic (MDP-level) and tactical (policy-level) reasoning. At the MDP level, we model cyberattacks as a Partially Observable Markov Decision Processes (POMDPs) to expose exploration-exploitation dynamics and phase-aware behavioural shifts. At the policy level, we analyse the temporal evolution of Q-values and use Prioritised Experience Replay (PER) to surface critical learning transitions and evolving action preferences. Evaluated across CyberBattleSim environments of increasing complexity, our framework offers interpretable insights into agent behaviour at scale. Unlike previous explainable RL methods, which are often post-hoc, domain-specific, or limited in depth, our approach is both agent- and environment-agnostic, supporting use cases ranging from red-team simulation to RL policy debugging. By transforming black-box learning into actionable behavioural intelligence, our framework enables both defenders and developers to better anticipate, analyse, and respond to autonomous cyber threats.
Keywords:
Explainable Artificial Intelligence Reinforcement Learning
Autonomous Cyber Operations Adversarial AI Cyberattack Modelling
1 Introduction
Modern enterprise networks face persistent threats from sophisticated adversaries who exploit misconfigurations to conduct lateral movement and privilege escalation attacks. As these threats grow in scale and complexity, static and reactive defences often fall short. They lack the agility to detect and respond to dynamic, multi-stage campaigns, highlighting the need for intelligent and adaptive security solutions [17]. Reinforcement Learning (RL)-based autonomous agents have emerged as promising solutions for modelling such adversarial behaviour, with growing use in platforms like CyberBattleSim [18], MITRE’s FARLAND [11], and CybORG [1].
Despite their increasing adoption, RL agents remain opaque black boxes, providing little visibility into why particular actions are taken, how strategies evolve, or which factors drive success or failure. This lack of interpretability poses barriers for both cybersecurity professionals and RL developers. For defenders, it limits the ability to anticipate threats or understanding adversarial reasoning. For developers, it hinders debugging, validation, and policy refinement in complex environments where decisions unfold over time under uncertainty.
Explainable Reinforcement Learning (XRL) aims to bridge this gap by converting RL agent behaviour into interpretable, context-aware insights that support trust, transparency, and safe deployment [22]. However, most prior explainability work in cybersecurity is limited to narrow tasks (e.g., malware detection, blockchain mining, autonomous driving, or small-scale intrusion detection [23, 6, 10, 2, 15]) and often relies on shallow, post-hoc methods. Few approaches support structured, temporally-aware explainability for RL agents operating in realistic, sequential, and partially observable adversarial settings [12, 9].
To address this gap, we introduce a unified, multi-layer explainability framework tailored for RL agents in cybersecurity. While we demonstrate it with autonomous attacker agents, the framework is agent- and environment-agnostic and applies equally to defenders. It provides two complementary lenses. At the MDP level, we model the environment as a Partially Observable Markov Decision Process (POMDP), enabling analysis of exploration-exploitation dynamics and behavioural shifts across early and late attack phases. At the policy level, we track the temporal Q-value evolution and incorporate Prioritised Experience Replay (PER) to identify critical learning transitions and preference formation over time. Together, these layers reveal how the agent behaviour matures, identifying intent, surfacing key decision points, and clarifying strategic inflection. This empowers analysts to trace agent learning over time and supports use cases ranging from red-team planning to model debugging.
We evaluate our framework using Microsoft’s CyberBattleSim [18], which simulates attacker behaviour in networked environments with privilege escalation, credential reuse, and lateral movement. Across environments of increasing complexity, our framework demonstrates how strategic patterns and tactical choices emerge under uncertainty. While our analysis focuses on attackers, the methodology generalises to defenders and other sequential decision-making domains as well such as malware containment, deception planning, or autonomous incident response. This positions our framework as a unified approach for explaining both red- and blue-team RL agents, enabling phase-aware defence design, training simulation enhancement, and policy transparency in autonomous cyber operations. Our key contributions are as follows:
  • Multi-layer explainability framework: We present a unified, multi-layer explainability framework for RL-based attacker agents. The framework integrates strategic (MDP-level) and tactical (policy-level) insights, bridging the transparency gap and enabling end-to-end interpretability of autonomous attacker behaviour.
  • MDP-level explainability: We model attacker behaviour as POMDP, enabling interpretable analysis of uncertainty handling and strategic progression through early and late phases of attack.
@@ 1 Introduction
  • Policy-level explainability: We use temporal Q-value tracking and prioritised experience replay to reveal how action preferences and critical decisions evolve during learning.
  • Empirical analysis: We conduct rigorous experiments in CyberBattleSim environments of increasing scale and complexity, showcasing the framework’s scalability and insightfulness.
  • Operational and developmental utility: Our framework supports both cyber defenders and RL developers by providing interpretable signals for anticipating behaviour, debugging policies, and designing adaptive responses.
2 Related Work
2.1 Reinforcement Learning for Cybersecurity
2.1.1 Simulation Environments.
@@ 2.1 Reinforcement Learning for Cybersecurity
2.1.2 RL-based Cyber Defence and Adversarial Simulation.
Recent work has explored both defensive and offensive RL agents. Thompson et al. [19] used entity-based RL with transformer policies to enable defensive generalisation in dynamic networks. Goel et al. [7] co-trained attacker and defender policies in a Stackelberg game for Active Directory protection. On the offensive side, Sultana et al. [16] trained deep RL agents to learn multi-layered cyberattack strategies across network, service, and application layers. Other approaches use RL to simulate red team behaviour for defender training [3, 5].
@@
2.2 Explainable Reinforcement Learning (XRL)
2.2.1 General XRL Techniques.
@@ 2.2 Explainable Reinforcement Learning (XRL)
Explainable reinforcement learning methods fall into two categories: intrinsic approaches [2], which embed interpretability during training (e.g., policy distillation, influence modelling), and post-hoc techniques [8, 10], which generate explanations after training using tools like SHAP, LIME, or counterfactual reasoning. For example, Alabdulkarim et al. [2] developed an intrinsic approach by training influence predictors alongside RL policies, enabling real-time analysis of how different actions shape agent behaviour over time. In contrast, Mathes et al. [10] introduced CODEX, a post-hoc, cluster-based XRL technique that semantically groups and summarizes both factual and counterfactual behaviours in environments such as MiniGrid and StarCraft II. Gyevnar et al. [8] proposed a social XAI framework for multi-agent autonomous driving, using post-hoc counterfactual causal selection to generate contrastive explanations for sequential decisions.
2.2.2 XRL Techniques in Cybersecurity.
@@
XRL in cybersecurity has largely targeted static classification tasks (e.g., malware detection [23], intrusion analysis [15]) rather than sequential decision-making. Yu et al. [23] incorporated explainability into malware mutation agents, while Sharma et al. [15] explored adversarial XRL for classifier sensitivity. Foley et al. [6] proposed a post-hoc XRL framework for defender agents in CybORG using SHAP, feature ablations, and state visualisations to improve situational awareness. However, their attacker agents follow fixed, hierarchical rules, limiting insight into real-time red-team adaptation.
@@
By contrast, our work places the attacker at the centre of the explanation. We introduce a general, environment-agnostic XRL framework that captures attacker behaviour across both strategic (MDP-level) and tactical (policy-level) layers. Unlike prior work, it supports both during- and post-training analysis and reveals how adversarial agents adapt over time. This empowers defenders to pre-empt adversarial tactics and developers to refine agent training, while offering actionable insight into sequential behaviours that remain opaque in existing approaches.
3 Problem Formulation
@@
Enterprise networks are increasingly targeted by lateral movement attacks, where adversaries navigate interconnected systems to escalate privileges and compromise critical assets. In Microsoft’s CyberBattleSim, such networks are modelled as directed graphs $G=(V,E)$, where nodes $v\in V$ represent machines (e.g., workstations, servers, domain controllers) and edges $e\in E$ represent communication links. Each node has an associated reward $r(v)$, reflecting its operational importance. An attacker, modelled as a RL agent, seeks to maximise cumulative reward $R=\sum_{v\in V_{C}}r(v)$, where $V_{C}\subseteq V$ denotes the set of compromised nodes. The agent operates under POMDP, reflecting the uncertainty and limited visibility typical of real-world cyberattacks. While RL agents can learn effective attack policies in this setting, their decision-making processes remain opaque. Traditional explainability methods, such as static feature attribution and post-hoc visualisations, fall short of capturing the sequential, adaptive nature of adversarial behaviour. As a result, defenders lack insight into when, why, and how attackers pivot laterally, escalate privileges, or shift tactics, insights that are critical for proactive defence and trust in autonomous systems.
4 CyberBattleSim Platform
@@
CyberBattleSim is an open-source simulation platform designed to model adversarial cyber operations within enterprise network environments. Networks are represented as directed graphs, where nodes encapsulate system configurations such as operating systems, services, known vulnerabilities, firewall rules, and edges represent connectivity. The environment is inherently attacker-centric: it trains RL agents to autonomously discover and exploit vulnerabilities through local privilege escalation, remote exploitation, and credential-based lateral movement. Unlike CybORG [1], which emphasises defender modelling and blue-team coordination, CyberBattleSim provides a focused setting for evaluating autonomous red-team strategies. Defender agents are not modelled explicitly; instead, defensive behaviour is encoded through static network configurations, access control rules, and hidden vulnerabilities.
@@
Agents begin with limited visibility and explore to uncover system properties and reachable nodes. The reward model incentivises strategic behaviour, assigning higher rewards to the compromise of mission-critical nodes (e.g., 100 for a domain controller vs. 10 for a workstation). Successfully breaching a node may also reveal previously hidden segments of the network, enabling progressive expansion of the attack agent’s surface. CyberBattleSim supports rich action spaces and complex environment dynamics, making it well-suited for studying both autonomous attacker behaviour and explainability in sequential, partially observable domains.
4.1 CyberBattleSim Environments

### v2

@@ Abstract
Reinforcement Learning (RL) agents are increasingly used to simulate sophisticated cyberattacks, but their decision-making processes remain opaque, hindering trust, debugging, and defensive preparedness. In high-stakes cybersecurity contexts, explainability is essential for understanding how adversarial strategies are formed and evolve over time. In this paper, we propose a unified, multi-layer explainability framework for RL-based attacker agents that reveals both strategic (Markov Decision Process (MDP)-level) and tactical (policy-level) reasoning. At the MDP-level, we model cyberattacks as a Partially Observable Markov Decision Process (POMDP) to expose exploration-exploitation dynamics and phase-aware behavioural shifts. At the policy-level, we analyse the temporal evolution of Q-values and use Prioritised Experience Replay (PER) to surface critical learning transitions and evolving action preferences. Evaluated across CyberBattleSim environments of increasing complexity, our framework offers interpretable insights into agent behaviour at scale. Unlike previous explainable RL methods, which are predominantly post-hoc, domain-specific, or limited in depth, our approach is both agent- and environment-agnostic, supporting use cases such as red-team simulation, RL policy debugging, phase-aware threat modelling and anticipatory defence planning. By transforming black-box learning into actionable behavioural intelligence, our framework enables both defenders and developers to better anticipate, analyse, and respond to autonomous cyber threats.
I Introduction
Mapping how attack paths emerge and evolve within a network is critical for understanding systemic vulnerabilities and anticipating adversarial behaviour[12]. To automate this process, recent work increasingly trains reinforcement learning (RL) agents in red-team simulators and cyber-range environments such as CyberBattleSim [17] and CybORG [7], where autonomous attacker agents learn to explore, pivot, reuse credentials, and escalate privileges through sequential interaction [25]. These RL agents can uncover complex, multi-step exploitation strategies that are difficult to enumerate manually, enabling systematic stress-testing of enterprise-scale defences [11, 13].
However, despite their effectiveness, the decision-making processes of RL-based attacker agents remain largely opaque. In practice, developers observe aggregate reward curves and final success rates, while the internal learning dynamics governing why, when, and how agents pivot, escalate, or persist, including policy evolution, exploration–exploitation transitions, and convergence or collapse, remain hidden. As a result, it is often unclear whether observed behaviour reflects coherent adversarial reasoning or degenerate learning patterns, such as repeatedly executing low-value actions (e.g., connect(GIT) or connect(PING)) that inflate reward without advancing the attack. This opacity undermines confidence in the realism and reliability of learned behaviours and prevents reliable debugging, benchmarking, and tuning of RL-driven adversaries for trustworthy red-team simulation.
Explainable Reinforcement Learning (XRL) seeks to address these challenges by making agent decision processes interpretable, clarifying why particular actions are selected, how policies evolve, and which environmental cues influence behaviour [29]. However, interpretability in RL poses fundamentally different challenges than in supervised learning. RL agents learn through continual interaction with uncertain, partially observable environments, producing non-stationary policies whose decision logic evolves throughout training [8]. As a result, most existing XRL methods remain fundamentally post-hoc, focusing on explaining fully trained policies and are typically evaluated in simplified or fully observable settings [14, 16, 15, 19].
Within cybersecurity, explainability efforts have further concentrated on narrow supervised tasks, such as malware detection, using feature-attribution techniques (e.g., SHAP, LIME) that provide limited insight into how sequential attacker strategies form, adapt, or fail over time [31, 9, 1, 23, 20, 2]. Recent behavioural audits of trained RL agents [6] offer informative summaries of final policies, yet fail to expose the training-time learning dynamics that generate, stabilise, or destabilise adversarial behaviour. This gap motivates the need for explainability mechanisms that operate during learning itself, rather than only after policy convergence.
Notably, many of the phenomena that determine behavioural realism and learning stability arise during training, not after convergence. In complex, partially observable cyber environments, agents may prematurely commit to narrow tactics, overfit early discoveries, or continue prioritising stale experiences long before such pathologies become visible in aggregate performance metrics. Training-time explainability is therefore not merely descriptive, but a practical diagnostic capability for identifying instability, reward misalignment, and premature exploitation early enough for intervention. Addressing this requires explainability across both high-level strategy formation and low-level policy optimisation.
To address this gap, we introduce a unified multi-layer explainability framework for RL-based cyber agents that makes policy formation observable as learning unfolds. Rather than providing static, post-hoc explanations of converged policies, our framework captures how strategies emerge, adapt, and stabilise during training, providing a principled diagnostic lens for analysing learning dynamics in adversarial environments.
The framework operates across two complementary levels.
1) At the strategic level (Markov Decision Process (MDP)), attacker–environment interaction is modelled as a partially observable Markov decision process (POMDP), enabling interpretable analysis of exploration–exploitation dynamics and phase-wise behavioural transitions.
@@ I Introduction
2) At the tactical level (policy), we analyse temporal Q-value evolution and Prioritised Experience Replay (PER) to identify learning inflection points and the consolidation of action preferences.
Together, these layers characterise how agent behaviour matures over time, revealing dominant strategies, key decision points, and adaptive shifts under uncertainty.
Attacker behaviour in our framework is not governed by fixed heuristics or predefined attack templates, but emerges directly from policy learning through environment interaction. Explainability is derived from training-time signals, including Q-value dynamics, temporal-difference (TD) error trajectories, policy-logit variance, and replay priorities, that expose when and why agents favour particular tactics (e.g., reconnaissance, lateral movement, or privilege escalation), how strategic preferences stabilise or collapse, and under what conditions novel behaviours arise. By grounding explanation in these signals, the framework enables early detection of instability, reward misalignment, and premature exploitation, well before such pathologies manifest in aggregate performance metrics. Unlike prior behavioural audits that characterise only final outcomes, our approach explicitly links internal learning dynamics to externally observable attack behaviour, exposing the causal mechanisms that govern policy formation under uncertainty.
We evaluate the proposed framework in Microsoft CyberBattleSim [17], which simulates attacker behaviour in networked environments involving privilege escalation, credential reuse, and lateral movement. Across environments of increasing scale and complexity, the framework reveals how strategic progression and tactical preferences develop throughout training, rather than only in final performance outcomes. Our primary experimental setup instruments a Deep Q-Learning (DQL) agent, a widely used baseline for attacker modelling in RL, where actions are naturally discrete and training proceeds episodically to capture bounded attack campaigns. To assess generality beyond value-based optimisation, we additionally evaluate the framework using the policy-gradient Proximal Policy Optimisation (PPO) algorithm. Collectively, these experiments demonstrate that training-time explainability exposes critical learning dynamics, including replay-induced instability, premature policy collapse, and value-propagation distortions, that remain fundamentally invisible under reward-only evaluation. We further show that naïvely applying prioritised experience replay (PER), commonly assumed to improve sample efficiency, can degrade learning in large cyber environments, a failure mode not captured by aggregate performance metrics. We further link policy-level collapse to internal confidence–uncertainty signals and externally observable behavioural lock-in.
While our evaluation focuses on attacker agents for red-team simulation, the proposed explainability primitives are algorithm- and environment-agnostic and extend naturally to other sequential cyber decision-makers, including defenders performing containment, deception, or adaptive response. Beyond analysis, the framework supports several practical capabilities for developing and validating autonomous cyber agents.
First, our framework enables targeted shaping of emergent behaviour. For instance, red teams or simulator developers may aim to induce specific adversarial styles, e.g., slow, stealthy reconnaissance reminiscent of APTs versus rapid smash-and-grab attacks. By analysing when agents shift from scanning to exploitation or develop preferences for high-impact actions, our framework provides visibility into whether training regimes are producing the intended behavioural profiles. This enables iterative tuning of exploration strategies, reward shaping, or environment complexity to guide learning toward more interpretable and mission-aligned attacker behaviour.
Second, it supports benchmarking and comparison of training configurations. Rather than relying solely on reward curves, our framework allows developers to track how strategy formation varies across runs. For example, one might ask: Does prioritised experience replay lead to earlier lateral movement? Does PER accelerate learning or instead introduce replay bias that destabilises convergence in large environments? Does a denser reward structure encourage premature escalation? Our multi-layer signals, capturing Q-value evolution, TD-error spikes, and exploration ratios, allow comparative audits of learning dynamics across different agent setups.
Third, it enables early diagnosis and control of learning dynamics. In complex, partially observable environments such as CyberBattleSim, RL agents may stall in early phases, fail to escalate, or converge on suboptimal behavioural patterns. Our framework exposes the underlying replay- and optimisation-driven mechanisms governing these dynamics, such as TD-error evolution, replay-priority shifts, entropy contraction, and dominant-action formation, well before such effects become visible in aggregate performance metrics. This enables earlier diagnosis, targeted intervention, and more reliable shaping of agent behaviour during training and simulation development.
Taken together, these capabilities position our proposed framework not only as an interpretability tool, but as a practical development aid for researchers and practitioners building reliable, explainable autonomous agents for cyber operations.
Our key contributions are as follows:
  • A training-time, multi-layer explainability framework for autonomous attacker agents, integrating strategic (MDP-level) and tactical (policy-level) analyses to expose how adversarial strategies emerge, stabilise, or collapse during learning.
  • Empirical insights into adversarial RL learning dynamics, including exploration–exploitation phase transitions, action-dominance formation, replay pathologies under PER, and policy saturation effects that are not observable from reward curves alone.
  • Cross-algorithm validation across value-based (DQL) and policy-gradient (PPO) agents, demonstrating that key behavioural patterns persist despite differing optimisation dynamics.
  • Practical diagnostic utility for red-team simulation developers, enabling debugging, benchmarking, and tuning of autonomous attacker agents to improve realism, stability, and controllability in cyber-range environments.
II Related Work
II-A Reinforcement Learning for Cybersecurity
II-A1 Simulation Environments
@@
II-A2 RL-based Cyber Defence and Adversarial Simulation
Recent work has explored both defensive and offensive RL agents. On the defensive front, Thompson et al. [26] used entity-based RL with transformer policies to enable generalisation in dynamic networks. Goel et al. [10] co-trained attacker and defender policies in a Stackelberg game for Active Directory protection. Wiebe et al. [30] employed cooperative MARL to enable distributed defenders to coordinate joint responses in complex host-based attack scenarios. On the offensive side, Sultana et al. [24] trained deep RL agents for multi-layered cyberattack strategies across network, service, and application layers. Other approaches simulate red team behaviour using RL to support defender training [3, 5].
@@ II-A2 RL-based Cyber Defence and Adversarial Simulation
II-B Explainable Reinforcement Learning (XRL)
II-B1 General XRL Techniques
@@
Explainable reinforcement learning methods fall into two categories: intrinsic approaches [1], which embed interpretability during training (e.g., policy distillation, influence modelling), and post-hoc techniques [14, 16], which generate explanations after training using tools like SHAP [2], LIME [20], or counterfactual reasoning. For example, Alabdulkarim et al. [1] developed an intrinsic approach by training influence predictors alongside RL policies, enabling real-time analysis of how different actions shape agent behaviour over time. In contrast, Mathes et al. [16] introduced CODEX, a post-hoc, cluster-based XRL technique that semantically groups and summarizes both factual and counterfactual behaviours in environments such as MiniGrid and StarCraft II. Gyevnar et al. [14] proposed a social XAI framework for multi-agent autonomous driving, using post-hoc counterfactual causal selection to generate contrastive explanations for sequential decisions. Madumal et al. [15] use structural causal models to generate post-hoc contrastive explanations for RL agents, enabling “why” and “why not” reasoning through counterfactuals, demonstrating their approach via human studies in StarCraft II. Nashed et al. [19] introduced a unified structural causal model framework for post-hoc explanation in sequential decision-making, supporting multiple explanation types via exact and approximate inference, and evaluating on abstract MDPs and user studies.
(a) Early stage: Initial foothold (Client node owned)
(b) Early stage: Website discovered
(c) Mid-stage: Expanded discovery and node compromise
(d) Mid-stage: Remote exploits and lateral movement
(e) Late-stage: Full visibility with partial control
(f) Late-stage: Full visibility and expanded control
Figure 1: Progressive discovery and control escalation by the RL-based attacker agent in the ToyCTF environment. Figure (a–b) In the early-stages, the agent initially controls only the Client node and discovers the adjacent Website node. Figure (c–d) In the mid-stage, the agent uncovers additional nodes and expands its control via remote exploits and lateral movements. Figure (e–f) In the late-stages, the agent attains full visibility and compromises additional critical nodes.
II-B2 XRL Techniques in Cybersecurity
@@
XRL in cybersecurity has largely targeted static classification tasks (e.g., malware detection [31], intrusion analysis [23]) rather than sequential decision-making. Yu et al. [31] incorporated explainability into malware mutation agents, while Sharma et al. [23] explored adversarial XRL for classifier sensitivity. For instance, they used adversarial perturbations to identify minimal input changes affecting classification outcomes, but such methods fail to capture the sequential, adaptive nature of cyber operations. Foley et al. [9] proposed a post-hoc XRL framework for defender agents in CybORG using SHAP, feature ablations, and state visualisations to improve situational awareness. However, their attacker agents follow fixed, hierarchical rules, limiting insight into real-time red-team adaptation.
@@
By contrast, our work centres the attacker in the explanatory process, introducing a general, environment-agnostic XRL framework that captures behaviour across both strategic (MDP-level) and tactical (policy-level) layers. Unlike existing approaches that are predominantly post-hoc, limited to static tasks, or tied to specific environments or agents, our framework supports both training-time and post-training analysis. This dual capability captures the temporal evolution of strategies as they emerge and adapt under uncertainty, enabling defenders to anticipate adversarial tactics and developers to refine agent training, while delivering actionable, interpretable insights that remain opaque in existing approaches.
III Problem Formulation
@@
Enterprise networks are increasingly targeted by lateral movement attacks, where adversaries navigate interconnected systems to escalate privileges and compromise critical assets. In Microsoft’s CyberBattleSim, such networks are modelled as directed graphs $G=(V,E)$, where nodes $v\in V$ represent machines (e.g., workstations, servers, domain controllers) and edges $e\in E$ represent communication links. Each node has an associated reward $r(v)$, reflecting its operational importance. An attacker, modelled as an RL agent, seeks to maximise cumulative reward $R=\sum_{v\in V_{C}}r(v)$, where $V_{C}\subseteq V$ denotes the set of compromised nodes. The agent operates under POMDP, reflecting the uncertainty and limited visibility typical of real-world cyberattacks. While RL agents can learn effective attack policies in this setting, their decision-making processes remain opaque. Traditional explainability methods, such as static feature attribution and post-hoc visualisations, fall short of capturing the sequential, adaptive nature of adversarial behaviour. As a result, defenders lack insight into when, why, and how attackers pivot laterally, escalate privileges, or shift tactics, insights that are critical for proactive defence and trust in autonomous systems.
IV CyberBattleSim Platform
@@
CyberBattleSim is an open-source simulation platform designed to model adversarial cyber operations within enterprise network environments. Networks are represented as directed graphs, where nodes encapsulate system configurations such as operating systems, services, known vulnerabilities, firewall rules, and edges represent connectivity. The environment is inherently attacker-centric: it trains RL agents to autonomously discover and exploit vulnerabilities through local privilege escalation, remote exploitation, and credential-based lateral movement. Unlike CybORG [7], which emphasises defender modelling and blue-team coordination, CyberBattleSim provides a focused setting for evaluating autonomous red-team strategies. CyberBattleSim does not model an explicit, adaptive defender agent; instead, defensive behaviour is encoded implicitly through static network configurations, access control rules, and vulnerability placement.
@@
Agents begin with limited visibility (Figure 1) and explore to uncover system properties and reachable nodes. The reward model incentivises strategic behaviour, assigning higher rewards to the compromise of mission-critical nodes (e.g., 100 for a domain controller vs. 10 for a workstation). Successfully breaching a node may also reveal previously hidden segments of the network, enabling progressive expansion of the attack agent’s surface. CyberBattleSim provides a structured yet non-trivial attacker action space and partially observable environment dynamics, making it suitable for studying how autonomous attacker behaviour and learning dynamics can be interpreted in sequential decision-making settings.
IV-A CyberBattleSim Environments